Are you supposed to use Order Of Operations when you Evaluate Expressions like:
(2+1-2+1)/2
When using the Order Of Operations,I get 0, and when not using them, I get 1.

Please help?

Answer by 303795(602)

Yes. You should be using the Order of Operations. (The actual terms used may differ depending where in the world you are and the abbreviations in use there. Here in Western Australia the abbreviation is BIMDAS)
The Order of Operations is
1 Brackets
2 Indices and powers
3 Multiplication and division
4 Addition and subtraction
If you have both multiplication and division then they are done from left to right. You do not do all the multiplications first then do the divisions.
If you have both addition and subtraction then they are done from left to right. You do not do all the additions first then do the subtractions.
In your example all the calculations inside the bracket need to be done first so working from left to right.
2+1=3
3-2=1
1+1=2
So the value inside the brackets simplifies to 2.
When this is divided by 2 the answer is 1.